OpenStudy (anonymous):

The ratio of all-star baseball cards to regular baseball cards in Tony's collection is five to seven. After Tony bought four more all-star baseball cards, his all star cards make up half of his new collection. How many cards did Tony have before his new purchase?

ganeshie8 (ganeshie8):

try seting up two equations and solve

ganeshie8 (ganeshie8):

say, all-star baseball cards = x regular baseball cards = y \(\large \frac{x}{y} = \frac{5}{7}\) ----------------------(1)

ganeshie8 (ganeshie8):

After Tony bought four more all-star baseball cards, his all star cards make up half of his new collection. \(\large x+4 = \frac{(x+4)+y}{2}\) ----------------(2)

ganeshie8 (ganeshie8):

two equations, two unknowns you can solve x and y and find 'x+y'

OpenStudy (anonymous):

is it 24? :o

ganeshie8 (ganeshie8):

correct x+y = 24 good work !!

OpenStudy (anonymous):

thanks :]

ganeshie8 (ganeshie8):

np :)